OpenStudy (anonymous):

What is the product of 6x+3 and 2x^2-5? Show all work.

OpenStudy (anonymous):

I dont see where it shows how to solve it or the answer.

OpenStudy (amistre64):

to solve it you simply multiply it together.....

OpenStudy (anonymous):

How do you do that?

OpenStudy (amistre64):

by applying the distributive property thru the problem

OpenStudy (amistre64):

8(6) = 48 (5+3)(2+4) = 48 (5+3)(2) + (5+3)(4) = 48 5(2) + 3(2) + 5(4) + 3(4) = 48 10 + 6 + 20 + 12 = 48 16 + 32 = 48 48 = 48 thats how we distribute....

OpenStudy (anonymous):

So the answer is 48?

OpenStudy (anonymous):

no he was just giving you an example . you have to do 4 multiplications

OpenStudy (anonymous):

\[(6x+3)(2x^2-5)\]

OpenStudy (anonymous):

\[6x\times 2x^2+6x\times -5+3\times 2x^2+3\times -5\]

OpenStudy (anonymous):

4 multiplications . then of course you actually do the multiplications as in \[6x\times 2x^2=12x^3\] \[6x\times -5=-30x\] \[3\times 2x^2=6x^2\] \[3\times -5=-15\]

OpenStudy (anonymous):

so your answer is \[12x^3-30x+6x^2-15\]